The atomic density of a solid is 5.85 ×10

28

 m

−3

×1028 m−3. Its electrical resistivity is 1.6 ×10

−8

Ωm

×10−8Ωm. Assume that each atom contributes one conduction electron:

2 points

The drift mobility (in m

2

 V

−1

 s

−1

m2 V−1 s−1) of the conduction electrons is:



1
Expert's answer
2021-09-22T07:06:12-0400
"\mu=\frac{1}{ne\rho}\\\mu=\frac{1}{(1.6\cdot10^{-8})(5.85\cdot10^{28})(1.6\cdot10^{-19})}=0.00668\frac{m^2}{VS}"